# Paper2Code: Automating Code Generation from Scientific Papers in Machine Learning

### v3

@@ 1 Introduction
Reproducibility lies at the heart of scientific progress, which enables researchers to validate findings, build upon prior work, and ultimately push the boundaries of knowledge [6, 3, 30]. However, across many disciplines, reproducing scientific results remains an enduring challenge. This is often due to incomplete documentation, missing experimental details, lack of access to data or proprietary tools, and, especially in machine learning research, the absence of corresponding code: for example, only average 19.5% of the papers accepted to top-tier machine learning conferences in 2024 provide their code implementations shown in Figure 1. As a result, researchers frequently invest substantial effort in reverse-engineering methods and experimental results from papers, a process that is both time-consuming and labor-intensive, subsequently slowing down the overall pace of scientific innovation.
Meanwhile, recent Large Language Models (LLMs) have shown outstanding capabilities in understanding and generating both natural language and programming code [10, 28, 35], with performances increasingly approaching or even surpassing that of domain experts in some scenarios. In addition, this progress has sparked growing interest in leveraging LLMs to accelerate scientific workflows, particularly in the early stages of ideation for new and valid research hypotheses [24, 47, 36, 2, 18, 48, 37]. Furthermore, some of these studies, as well as others focusing on later stages of automating experimental validations and improvements [14, 51, 4, 41], demonstrate the potential of LLMs to generate code and even carry out experiments end-to-end; however, they typically assume and heavily rely on access to pre-existing implementations, partial code snippets, or well-defined APIs. As such, it remains questionable whether generating complete and faithful implementations solely from research papers (without access to prior code, APIs, or additional supplementary materials) can be achievable.
@@
Figure 1: (a) PaperCoder overview. The proposed PaperCoder aims to transform given scientific papers (in machine learning domains) into code repositories, which consists of three sequential steps: planning, analysis, and coding. (b) Code availability, where blue bars indicate the total number of accepted papers and orange regions show those with officially released code (See Appendix B.1 for details on the availability calculation).
To validate the effectiveness of PaperCoder, we conduct extensive evaluations on a subset of recent machine learning papers accepted at top-tier venues in 2024, including ICLR, ICML, and NeurIPS referred to as our proposed Paper2Code benchmark (in short, Paper2CodeBench). Also, we incorporate the recent PaperBench benchmark [38] in our evaluation suite, enabling fine-grained evaluations of code implementations with some papers from ICML 2024. Then, on a battery of tests conducted not only with automated model-based evaluations (covering both reference-free and reference-based settings, conditional on the availability of author-released ground-truth repositories) but also with expert human evaluations (based on authors of original papers), PaperCoder demonstrates substantial improvements over baselines, generating more valid and faithful code repositories that could meaningfully support human researchers in understanding and reproducing prior work. Specifically, 88% of the generated repositories by PaperCoder are rated as the best over baselines, and 92% of human judges report that the generated repositories are indeed helpful. Also, further analyses show that each component of PaperCoder (consisting of planning, analysis, and generation) contributes to the performance gains, but also that the generated codebases can be executed, sometimes with only minor modifications (averaging 0.81% of total code lines) in cases where execution errors occur.
@@ 2 Related Work
LLMs have shown impressive capabilities in text understanding and generation [28, 10, 35], which have been widely utilized within specialized domains (beyond general tasks), such as mathematics, science, and coding [32, 43, 40]. Particularly, code-specialized LLMs [9, 8, 15] have received significant attention thanks to remarkable performance on various software engineering tasks [45], including software design and development [34, 13], requirements elicitation [27], and formal specification generation [25]. Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of (code-specialized) LLMs.
@@
Early work on LLM-based code generation typically focuses on single-file tasks, whose objective is to generate relatively short code snippets to solve isolated tasks, such as (algorithmic-level) programming competition problems or toy-level programming queries [19, 5, 1, 12]. However, as LLMs have advanced in comprehending and generating code with the ability to perform long-context reasoning, recent studies have increasingly shifted their attention toward more challenging repository-level coding tasks, which involve generating multi-file repositories that jointly account for architectural design, modular structure, and inter-file dependencies [21, 16, 39]. In particular, several recent efforts explore this emerging paradigm [50, 29], often adopting multi-agent or role-based frameworks to emulate realistic development workflows. For instance, ChatDev instantiates LLMs into role-playing agents that collaborate through structured dialogues [34], while MetaGPT implements a waterfall-style development pipeline with specialized agents for handling different stages of software engineering [13]. Beyond prior work, we explore the underexplored yet challenging task of transforming full-length, complex papers into repository-level implementations.
@@
LLMs have been increasingly adopted to support every stage of the scientific process from ideation to experimental validation [31, 2, 18, 48, 33, 7, 20, 44]; thereby, helping researchers overcome existing challenges and ultimately accelerate scientific discovery [17, 24, 47]. Specifically, in fields such as computer science and machine learning (where code-based experimentation is central), LLMs have been used to design, refine, and extend code implementations. Yet, many recent efforts in this space assume access to and build on top of the original codebase [14, 46, 4, 41], which significantly limits their applicability in real-world scenarios since such implementations are oftentimes unavailable (Figure 1). To address this, concurrent to our work, Starace et al. [38] introduces a benchmark dataset called PaperBench, evaluating the capability of existing agentic AI systems in reproducing papers with fine-grained metrics. Notably, on top of PaperBench (which emphasizes evaluation), we further complement and extend this line by focusing on methodological aspects of how to transform scientific papers into repository-level code implementations.
@@ 3 Method
3.1 Repository-Level Code Generation from Scientific Papers in Machine Learning
The goal of our repository-level code generation task is to automatically produce a code repository that faithfully implements the methods and experiments described in machine learning papers (especially for cases where authors do not release their code), to support reproducibility and accelerate scientific progress [26, 30]. Formally, we define this task as a function (or a model) $M$ that maps a research paper $R$ to a corresponding code repository $C$, as follows: $M(R)=C$. Here, $C$ is composed of multiple interdependent files $\{c_{1},c_{2},...,c_{n}\}$, each responsible for implementing different components of the methods and experiments described in $R$, but together they should form a cohesive pipeline.
It is worth noting that the most straightforward approach to instantiating $M$ is to instruct the LLM to generate the entire code repository, conditioned on the given paper, as follows: $M(R):=\texttt{LLM}(\mathcal{T}(R))$, where $\mathcal{T}$ is the prompt template that specifies the intended behavior of the LLM for the target task (including task descriptions, detailed instructions, and any other relevant context). Yet, generating a complete, modular, and faithful repository (spanning multiple files) in a single pass is extremely challenging, even for powerful LLMs, due to the inherent complexity of scientific papers and their corresponding implementations, the long-context limitations of current models, and the difficulty in maintaining consistent global structure and cross-file dependencies. Therefore, we propose a multi-stage and multi-agent framework that decomposes the overall task into smaller, manageable subtasks, each handled by a specialized agent tailored to a specific aspect of paper-to-code transformation.
@@ 3.1 Repository-Level Code Generation from Scientific Papers in Machine Learning
3.2 PaperCoder: LLM-Powered Multi-Agent Framework for Paper-to-Code Transformation
@@ 3.2.1 Planning
It is worth noting that, in contrast to implementation specifications designed explicitly for software development, scientific papers are written to communicate ideas and findings to humans. As a result, they often contain high-level motivations, persuasive narratives, and auxiliary details that are crucial for human understanding but noisy, loosely specified, or ambiguous from a software engineering perspective. To mitigate this challenge, we introduce a structured planning phase that transforms unstructured textual content into implementation-level abstractions. Also, we decompose the planning process into four sequential subcomponents (to simplify the task and reduce cognitive and computational load of LLM-powered agents at each step): 1) overall plan, 2) architecture design, 3) logic design, and 4) configuration generation. Formally, we define this as: $M_{\text{plan}}(R)\rightarrow P=\{o,d,l,g\}$, where $o$ is the overall plan, $d$ is the architecture design, $l$ is the logic design, and $g$ is the configuration file, with each stage using the outputs of the previous ones as contextual input. We then describe how each subcomponent is instantiated through specific objectives and instructions below.
@@
While the previous architecture design focuses on what to build, the logic design phase specifies how these components should be instantiated in practice by considering their dependencies in terms of overall execution flow. In other words, its primary goal is to translate abstract component-level relationships from the architecture design into an actionable execution order for the entire implementation. This step is crucial because individual modules often depend on shared utilities, configurations, or data loaders that are defined in other parts of the repository, and without an explicitly defined execution order, the code generation can result in failure or inconsistency (e.g., generating file B before file A when B imports modules from A). To address this, the logic design stage not only produces an ordered file list that dictates the sequence in which the files should be implemented and executed, but also further elaborates on the logic within each file; thereby, providing more fine-grained specifications. Formally, $M_{\text{plan}}^{(3)}(R,o,d):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(3)}(R,o,d))\rightarrow l$.
@@
In the last stage of planning, PaperCoder synthesizes a configuration file (namely, config.yaml) that includes key hyperparameters, model settings, and other runtime options based on prior outputs alongside the given paper. We note that, in addition to grounding the code generation process with the explicit configuration details, it enables researchers to easily review and adjust experimental configurations without modifying the source code. Formally, $M_{\text{plan}}^{(4)}(R,o,d,l):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(4)}(R,o,d,l))\rightarrow g$. We provide prompts used to elicit each planning output in Appendix E.
@@ 3.2.3 Coding
The final stage of our PaperCoder framework is the coding phase, where the complete code repository is produced. In particular, each file is generated based on all the available contextual information accumulated from the previous stages, including the overall plan, architecture design, logic design, configuration file, and file-specific analyses, as well as the original paper. Additionally, to ensure consistency across different files, we generate them sequentially according to the execution order (i.e., the ordered file list determined during the logic design stage). To be formal, the overall code generation process is defined as follows: for each file $f_{i}$, the corresponding code $c_{i}$ is generated as $M_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}):=\texttt{LLM}(\mathcal{T}_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}))\rightarrow c_{i}$, resulting in the complete code repository $C=\{c_{i}\}_{i=1}^{n=|F|}$. We note that this iterative formulation can ensure that $i$-th code is generated with full awareness of its dependencies and the evolving state of the repository.
4 Experimental Setup
We now describe the experimental setup, including datasets, methods, and evaluation metrics.
4.1 Datasets
@@
To validate the efficacy of our PaperCoder framework in transforming machine learning papers into their corresponding repository-level codes, we construct a new benchmark dataset, which we refer to as Paper2CodeBench. Specifically, we collect the accepted papers from recent machine learning venues (such as ICLR, ICML, and NeurIPS 2024) with the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, and filter them based on the availability of code with its total number of tokens less than 70,000, to ensure the full repository remains within reasonable processing limits of modern LLMs for generation and evaluation. Also, to maintain the benchmark quality, we perform model-based evaluation [22] with GPT-4o on all the collected repositories and select the top 30 from each venue, resulting in a total of 90 papers listed in Tables 1112, and 13. Also, we additionally consider 21 papers for human evaluation (See Table 15). In addition to Paper2CodeBench, we also use the recently released PaperBench Code-Dev [38], which consists of 20 papers from ICML 2024 with paper-specific rubrics annotated by humans. In particular, those rubrics are used to judge the correct implementation based on LLM-based evaluation.
4.2 Baselines and Our Model
@@
We target the novel problem of Paper2Code, and there are no baselines designed for it to enable direct comparison. Nevertheless, we consider several related approaches proposed to implement repository-level code (or the entire software) from natural language inputs (such as software requirements), in addition to the ablated variants of our full PaperCoder framework, as follows: ChatDev [34] is a multi-agent framework for software development, where several role-specific LLM-powered agents collaborate via structured dialogues; MetaGPT [13] similarity adopts a role-based multi-agent paradigm, but its development process is organized by the principle of Standardized Operating Procedures (SOPs); Abstract is a naive variant of our PaperCoder, which uses only the paper abstract for implementation; Paper, while using the full paper, prompts LLMs only once for code generation; PaperCoder (Ours) is our full framework, structured into three stages of planning, analysis, and code generation. In addition to them, for the PaperBench Code-Dev, we consider baselines suggested by it: Basic Agent is the agentic architecture that can run a predefined set of tools with the ReAct-style approach [49], built upon the basic agent from Inspect AI [^2] https://inspect.ai-safety-institute.org.uk/agents.html#sec-basic-agent, and Iterative Agent that extends Basic Agent, iteratively instructing the model to complete the next subtask.
4.3 Evaluation Setup
Recall that, as shown in Figure 1, the official code implementations of many papers are not available; however, manually annotating their corresponding code implementations to evaluate the quality of automatically generated code repositories is highly labor-intensive and challenging. To address this and ultimately perform the evaluation at scale, we design two model-based evaluation protocols: reference-based (when ground-truth code is available) and reference-free (when it is not), following the recent trends in using LLMs to judge the quality of model outputs [52, 11, 22], especially when they are long. In addition to this, we also perform human evaluations with the authors of the original papers, to ensure reliable judgments and to assess the quality of our model-based evaluations by measuring their correlation with human scores. We discuss each evaluation protocol in detail below.
Reference-Based Evaluation
In this evaluation setup, we use the official author-released repository as the gold standard if it is available, since it most accurately reflects the implementations intended by the authors, including the components they consider essential for reproducing the main results presented in the paper. Specifically, we prompt the evaluation model (such as o3-mini-high [^3] Unless otherwise stated, we use o3-mini-high due to strong code understanding and reasoning capability.) to judge the quality of the generated repository with respect to the gold repository, alongside the input paper as context (See Appendix E for the detailed prompt). During evaluation, the model then identifies key components (to be implemented), categorizes them into three severity levels (high, medium, and low), and critiques how well each component is implemented. After that, it returns the overall score on a 5-point Likert scale. We note that, to ensure the reliability of the model-based evaluation, we sample the outputs multiple times (for example, 8 generations per evaluation) and report the averaged score.
Reference-Free Evaluation
For cases where the official author-released code is not available, we introduce the reference-free evaluation protocol that leverages only the paper to assess the quality of its generated repository. In other words, similar to the reference-based evaluation protocol, the evaluation model is prompted to identify key components, categorize them by severity, and critique their implementations in the generated code, but they are performed solely based on the information provided in the paper without access to the ground-truth repository. The rest of the evaluation process, such as sampling and score averaging, follows the same setup as in the reference-based evaluation.
Human Evaluation
While model-based evaluation offers a scalable and automated way of assessment, we also conduct human evaluations to validate our PaperCoder based on the expert-grounded evaluation. Specifically, to ensure informed and accurate judgment, each participant is assigned a paper for which they are the first author. Furthermore, they are presented with multiple implementations generated by different approaches, and asked to rank them. We provide more details in Appendix A.2.
Evaluation on PaperBench Code-Dev
For evaluation on PaperBench Code-Dev [38], we follow their evaluation setup, measuring the score over the paper-specific rubrics with LLM-based evaluation.
@@ 4.3 Evaluation Setup
5 Experimental Results and Analysis

### v4

@@ 1 Introduction
Reproducibility lies at the heart of scientific progress, which enables researchers to validate findings, build upon prior work, and ultimately push the boundaries of knowledge (Collaboration, 2015; Baker, 2016; Pineau et al., 2021). However, reproducing scientific results remains an enduring challenge. This is often due to incomplete documentation, missing experimental details, lack of access to data or proprietary tools, and, especially in machine learning research, the absence of corresponding code: for example, only average 19.5% of the papers accepted to top-tier machine learning conferences in 2024 provide their code implementations shown in Figure 1. As a result, researchers frequently invest substantial effort in reverse-engineering methods and experimental results from papers, a process that is both time-consuming and labor-intensive, subsequently slowing down the overall pace of science.
Meanwhile, recent Large Language Models (LLMs) have shown outstanding capabilities in understanding and generating both natural language and programming code (Dubey et al., 2024; OpenAI, 2024; Reid et al., 2024), with performances increasingly approaching or even surpassing that of domain experts in some scenarios. In addition, this progress has sparked growing interest in leveraging LLMs to accelerate scientific workflows, particularly in the early stages of ideation for new and valid research hypotheses (Lu et al., 2024; Li et al., 2024; Yang et al., 2024; Si et al., 2024; Yamada et al., 2025; Schmidgall et al., 2025; Baek et al., 2025). Furthermore, some of these studies, as well as others focusing on later stages of automating experimental validations and improvements (Huang et al., 2024; Zhang et al., 2024; Trirat et al., 2024; Chan et al., 2025), demonstrate the potential of LLMs to generate code and even carry out experiments end-to-end; however, they typically assume and heavily rely on access to pre-existing implementations, partial code snippets, or well-defined APIs. As such, it remains questionable whether generating faithful implementations solely from papers (without access to prior code, APIs, or additional supplementary materials) can be achievable.
@@
Figure 1: (a) PaperCoder, which aims to transform given scientific papers into code repositories, consisting of planning, analysis, and coding steps. (b) Code availability, where blue bars indicate the total number of accepted papers and orange regions show those with officially released code (See Appendix B.1 for calculation details).
To validate the effectiveness of PaperCoder, we conduct extensive evaluations on a subset of recent machine learning papers from ICLR, ICML, and NeurIPS referred to as our proposed Paper2Code benchmark (in short, Paper2CodeBench). Also, we incorporate the recent benchmark (Starace et al., 2025) in our evaluation suite, enabling fine-grained evaluations of code implementations. Then, on a battery of tests conducted not only with automated model-based evaluations (covering both reference-free and reference-based settings, conditional on the availability of author-released ground-truth repositories) but also with expert human evaluations (based on authors of original papers), PaperCoder demonstrates substantial improvements over baselines, generating more valid and faithful code repositories that could meaningfully support human researchers in reproducing prior work. Specifically, 88% of the generated repositories by PaperCoder are rated as the best over baselines, and 92% of human judges report that the generated repositories are indeed helpful. Also, analyses show that each component of PaperCoder (consisting of planning, analysis, and generation) contributes to the performance gains, but also that the generated codebases can be executed, sometimes with only minor modifications (averaging 0.81% of total code lines) in cases where execution errors occur.
@@ 2 Related Work
LLMs have shown impressive capabilities in text understanding and generation (OpenAI, 2024; Dubey et al., 2024; Reid et al., 2024) and widely utilized for specialized domains (beyond general tasks), such as mathematics, science, and coding (Prabhakar et al., 2025; Wang et al., 2024; Trinh et al., 2024). Particularly, code-specialized LLMs (Hui et al., 2024; DeepSeek-AI et al., 2024; 2025) have received significant attention thanks to remarkable performance on various software engineering tasks (Xia et al., 2024), including software design and development (Qian et al., 2024; Hong et al., 2024), requirements elicitation (Mu et al., 2023), and formal specification generation (Luo et al., 2024). Our work aligns closely with this line of research, exploring and expanding upon the capabilities and applications of (code-specialized) LLMs.
@@
Early work on code generation typically focuses on single-file tasks, whose objective is to generate short code snippets to solve isolated tasks, such as (algorithmic-level) programming competition problems (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Li et al., 2022). However, as LLMs have advanced in comprehending and generating code with the long-context reasoning ability, recent studies have increasingly shifted their attention toward more challenging repository-level coding tasks, which involve generating multi-file repositories that jointly account for architectural design, modular structure, and inter-file dependencies (Liu et al., 2024; Jain et al., 2024; Tang et al., 2024). In particular, several recent efforts explore this emerging paradigm (Zhang et al., 2023; Ouyang et al., 2025), adopting multi-agent or role-based frameworks to emulate realistic development workflows. For instance, ChatDev instantiates LLMs into role-playing agents that collaborate through structured dialogues (Qian et al., 2024), while MetaGPT implements a waterfall-style development pipeline with specialized agents (Hong et al., 2024). Beyond prior work, we explore the underexplored task of transforming full, complex papers into repository-level code.
@@
LLMs have been adopted to support the scientific process from ideation to experimental validation (Popper, 1959; Qi et al., 2023; Li et al., 2024; Yang et al., 2024; D’Arcy et al., 2024; Liang et al., 2024; Baek et al., 2025; Weng et al., 2025); thereby, helping researchers overcome existing challenges and ultimately accelerate scientific discovery (Lehr et al., 2024; Lu et al., 2024; Yamada et al., 2025). Specifically, in fields such as computer science (where code-based experimentation is central), LLMs have been used to design, refine, and extend code implementations. However, many recent efforts in this space assume access to and build on top of the original codebase (Huang et al., 2024; Trirat et al., 2024; Xiang et al., 2025; Chan et al., 2025), which significantly limits their applicability in real-world scenarios since such implementations are oftentimes unavailable (See Figure 1). To address this, concurrent to our work, Starace et al. (2025) introduces a benchmark dataset called PaperBench, evaluating the capability of existing agentic AI systems in reproducing papers with fine-grained metrics. Notably, on top of PaperBench (which emphasizes evaluation), we further complement and extend this line by focusing on methodological aspects of how to transform scientific papers into repository-level code implementations.
@@ 3 Method
3.1 Repository-Level Code Generation from Machine Learning Papers
The goal of our repository-level code generation task is to automatically produce a repository that faithfully implements methods and experiments described in machine learning papers (especially for cases where authors do not release their code), to support reproducibility and accelerate scientific progress (Pineau et al., 2021; Magnusson et al., 2023). Formally, we define this task as a function (or a model) $M$ that maps a paper $R$ to a corresponding code repository $C$, as follows: $M(R)=C$. Here, $C$ is composed of multiple files $\{c_{1},c_{2},...,c_{n}\}$, each responsible for implementing different components of the methods and experiments in $R$, but together they should form a cohesive pipeline.
The most straightforward approach to instantiating $M$ is to instruct the LLM to generate the entire code repository, conditioned on the given paper, as follows: $M(R):=\texttt{LLM}(\mathcal{T}(R))$, where $\mathcal{T}$ is the prompt template that specifies the intended behavior of the LLM for the target task (including task descriptions, detailed instructions, and any other relevant context). Yet, generating a complete, modular, and faithful repository in a single pass is extremely challenging, even for powerful LLMs, due to the inherent complexity of scientific papers and their corresponding implementations, the long-context limitations of current models, and the difficulty in maintaining consistent global structure and cross-file dependencies. Therefore, we propose to decompose the overall task into smaller subtasks, each handled by a specialized agent tailored to a specific aspect of paper-to-code transformation.
@@ 3.1 Repository-Level Code Generation from Machine Learning Papers
3.2 PaperCoder: LLM-Powered Multi-Agent Framework for Paper-to-Code
@@ 3.2.1 Planning
It is worth noting that, in contrast to implementation specifications designed explicitly for software development, papers are written to communicate ideas and findings to humans. As a result, they often contain high-level motivations, persuasive narratives, and auxiliary details that are crucial for human understanding but noisy, loosely specified, or ambiguous from a software engineering perspective. To mitigate this, we introduce a planning phase that transforms unstructured textual content into implementation-level abstractions. Also, we decompose the planning process into four sequential subcomponents (to simplify the task and reduce cognitive load of LLM-powered agents at each step): 1) overall plan, 2) architecture design, 3) logic design, and 4) configuration generation. Formally, we define this as: $M_{\text{plan}}(R)\rightarrow P=\{o,d,l,g\}$, where $o$ is the overall plan, $d$ is the architecture design, $l$ is the logic design, and $g$ is the configuration file, with each stage using the outputs of the previous ones as contextual input. We then describe how each subcomponent is instantiated below.
@@
While the previous architecture design focuses on what to build, the logic design phase specifies how these components should be instantiated in practice by considering their dependencies in terms of overall execution flow. This step is crucial because individual modules often depend on shared utilities, configurations, or data loaders that are defined in other parts of the repository, and without an explicitly defined execution order, the code generation can result in failure or inconsistency (e.g., generating file B before file A when B imports modules from A). To address this, the logic design stage not only produces an ordered file list that dictates the sequence in which the files should be implemented and executed, but also further elaborates on the logic within each file; thereby, providing more fine-grained specifications. Formally, $M_{\text{plan}}^{(3)}(R,o,d):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(3)}(R,o,d))\rightarrow l$.
@@
In the last stage of planning, PaperCoder synthesizes a configuration file (config.yaml) that includes key hyperparameters, model settings, and other runtime options based on prior outputs alongside the given paper. We note that, in addition to grounding the code generation process with the explicit configuration details, it enables researchers to easily review and adjust experimental configurations without modifying the source code. Formally, $M_{\text{plan}}^{(4)}(R,o,d,l):=\texttt{LLM}(\mathcal{T}_{\text{plan}}^{(4)}(R,o,d,l))\rightarrow g$. We provide prompts used to elicit each planning output in Appendix D.
@@ 3.2.3 Coding
The final stage is the coding phase, where the complete code repository is produced. In particular, each file is generated based on all the available contextual information accumulated from the previous stages, including the overall plan, architecture design, logic design, configuration file, and file-specific analyses, as well as the original paper. Additionally, to ensure consistency across different files, we generate them sequentially according to the execution order (i.e., the ordered file list determined during the logic design stage). To be formal, for each file $f_{i}$, the corresponding code $c_{i}$ is generated as follows: $M_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}):=\texttt{LLM}(\mathcal{T}_{\text{code}}(R,P,f_{i},a_{i},\{c_{1},...,c_{i-1}\}))\rightarrow c_{i}$, resulting in the complete code repository $C=\{c_{i}\}_{i=1}^{n=|F|}$. We note that this iterative formulation can ensure that $i$-th code is generated with full awareness of its dependencies and the evolving state of the repository.
4 Experiment
We now describe the experimental setup and the experimental results with reproducibility analyses.
4.1 Experimental Setup
Datasets
@@ 4.1 Experimental Setup
To evaluate our PaperCoder, we construct a new benchmark (Paper2CodeBench). Specifically, we collect the accepted papers from recent machine learning venues (such as ICLR, ICML, and NeurIPS 2024) with the OpenReview API [^1] https://docs.openreview.net/reference/api-v2, and filter them based on the availability of code with its total number of tokens less than 70,000, to ensure the full repository remains within reasonable processing limits of modern LLMs for generation and evaluation. Also, to maintain the quality, we perform model-based evaluation (Liu et al., 2023) with GPT-4o on all the collected repositories and select the top 30 from each venue, resulting in a total of 90 papers listed in Tables 1718, and 19. Moreover, we additionally consider 21 papers for human evaluation (See Table 20). In addition to Paper2CodeBench, we also use the recently released PaperBench Code-Dev (Starace et al., 2025), which consists of 20 papers from ICML 2024 with paper-specific rubrics annotated by humans. In particular, those rubrics are used to judge the correct implementation based on LLM-based evaluation.
Baselines and Our Model
@@
We target the novel problem of Paper2Code, and there are no baselines designed for it to enable direct comparison. Nevertheless, we consider several related approaches proposed to implement repository-level code (or the entire software) from natural language inputs (such as software requirements), in addition to the ablated variants of our full PaperCoder framework, as follows: ChatDev (Qian et al., 2024) is a multi-agent framework for software development, where several role-specific LLM-powered agents collaborate via structured dialogues; MetaGPT (Hong et al., 2024) similarity adopts a role-based multi-agent paradigm, but its process is organized by the principle of Standardized Operating Procedures (SOPs); Abstract is a variant of our PaperCoder, which uses only the paper abstract for implementation; Paper, while using the full paper, performs one-shot code generation; PaperCoder (Ours) is our full framework, structured into three stages of planning, analysis, and code generation. Additionally, for the PaperBench Code-Dev, we consider baselines suggested by it: Basic Agent is the agentic architecture that can run a predefined set of tools with the ReAct-style approach (Yao et al., 2023), built upon the agent from Inspect AI [^2] https://inspect.ai-safety-institute.org.uk/agents.html#sec-basic-agent, and Iterative Agent that extends Basic Agent, iteratively instructing the model to complete the next subtask.
@@
Evaluation Setup
Recall that, as shown in Figure 1, the official code implementations of many papers are not available; however, manually annotating their corresponding code implementations to evaluate the quality of automatically generated code repositories is highly labor-intensive and challenging. To address this and ultimately perform the evaluation at scale, we design two evaluation protocols: reference-based (when ground-truth code is available) and reference-free (when it is not), following the recent trends in using LLMs as a judge (Zheng et al., 2023; Fu et al., 2024; Liu et al., 2023). In addition to this, we also perform human evaluations with the authors of the original papers, to ensure reliable judgments and to assess the quality of our model-based evaluations by measuring their correlation with human scores. We discuss each evaluation protocol in detail below.
  • Reference-Based Evaluation. We use the official author-released repository as the gold standard only if it is available, since it most accurately reflects the implementations intended by the authors, including the components they consider essential to their main ideas. Specifically, we prompt the model (such as o3-mini-high [^3] Unless otherwise stated, we use o3-mini-high due to strong code understanding and reasoning capability.) to judge the quality of the generated repository with respect to the gold repository, alongside the input paper as context (See Appendix D for the detailed prompt). The model then identifies components (to be implemented), categorizes them into three severity levels (high, medium, and low), and critiques how well each component is implemented. After that, it returns the overall score on a 5-point Likert scale. We note that, to ensure the reliability of the model-based evaluation, we sample multiple outputs (e.g., 8) and report the average score.
  • Reference-Free Evaluation. For cases where the official author-released code is not available, we introduce the reference-free evaluation protocol that leverages only the paper to assess the quality of its generated repository. Similar to the reference-based evaluation, the evaluation model is prompted to identify key components, categorize them by severity, and critique their implementations in the generated code, but they are performed solely based on the information provided in the paper. The rest of the evaluation process, such as sampling and score averaging, follows the same setup.
  • Human Evaluation. While model-based evaluation offers a scalable and automated way of assessment, we also conduct human evaluations to validate our PaperCoder based on expert-grounded evaluation. Specifically, to ensure informed and accurate judgment, each participant is assigned a paper for which they are the first author. Also, they are presented with multiple implementations generated by different approaches, and asked to rank them. We offer more details in Appendix A.2.
Lastly, for evaluation on the PaperBench Code-Dev benchmark (Starace et al., 2025), we follow their evaluation setup, measuring the score over the paper-specific rubrics with LLM-based evaluation.
4.2 Experimental Results and Analysis